OpenStudy (anonymous):

Solve this system of linear equations. 6x+6y=0 -10x+4y=28

OpenStudy (anonymous):

divide first equation by 6 (both sides) and second - by 2 you'll have: x+y=0 -5x+2y=14 right?

OpenStudy (anonymous):

from the first one you can get: y=-x substitute y for (-x) in second: -5x +2*(-x)=14 solve it for x

OpenStudy (anonymous):

-7x=14 x=-2 from the first equation we know that y=-x so, y=2 the answer: x=-2, y=2 or (-2,2)

OpenStudy (anonymous):

The answer? (:

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

see above... x=-2, y=2 or (-2,2)